# From Sub-Ability Diagnosis to Human-Aligned Generation: Bridging the Gap for Text Length Control via MarkerGen

### v2

@@ 1 Introduction
However, the ongoing enhancements in LLM capabilities have yet to deliver the expected performance in LCTG while ensuring semantic integrity Foster et al. (2024); Wang et al. (2024); Song et al. (2024). Yuan et al. (2024) reports that even advanced LLMs (e.g., GPT-4 Turbo (OpenAI, 2023)) violate the given length constraints almost 50% of the time. To address this, training-based methods (Park et al., 2024; Yuan et al., 2024; Jie et al., 2023; Li et al., 2024b) have been studied to reinforce LLMs’ adherence to length constraints, yet they face two key challenges: (1) Limited generalization: Since text types are diverse and length constraints vary widely (e.g., ranging from an exact 500 words to coarse intervals like 500–600 words or below 500 words), training-based methods often fail to generalize effectively across different settings, as demonstrated in Appendix E.1. (2) Inferior controllability: These methods strengthen LCTG by enforcing implicit length modeling during generation in a top-down manner via training, lacking the decomposition and targeted enhancement of LCTG sub-capabilities, thereby limiting their progress (Retkowski and Waibel, 2024).
@@ 2.5 LCTG Error
LCTG error is primarily attributed to fundamental deficiencies in length modeling, following the order of Counting Error $>$ Perception Error $>$ Aligning Error $\gg$ Planning Error. Thus, as counting interval increases, the accumulation of counting errors leads to a corresponding rise in LCTG error.
@@ 3.2 Three-Stage Decoupled Generation
See Appendix 2 for prompts of each stage.
@@ 4.3 Generalizability
Table 3 also shows MarkerGen’s strong performance across benchmarks with varying length scale (18–1450). To further investigate, we analyze progressively increasing the target length from 100 to 400. The results in Table 4 show a declining trend in MarkerGen’s error rate, which can be attributed to the auxiliary marker insertion decoding mechanism that prevents error accumulation from implicit modeling.
@@
In addition to exact length constraints, users may impose range-based limits. We evaluate $E_{r}$, the proportion of responses violating these constraints. Table 4 shows that MarkerGen maintains an $E_{r}$ below 3% in all cases, significantly lower than the baseline.
@@ 4.4 Ablation Studies
The introduction of explicit length markers in the two-stage scheme leads to a substantial reduction in LCTG error relative to the implicit baseline (8.7 → 2.66). However, this scheme places greater emphasis on length modeling, which consequently diminishes text quality (4.45 → 4.28). In comparison, the three-stage scheme achieves a better balance by decoupling semantic and length modeling, thereby improving both length control and text quality.
@@ Appendix D Detailed Benchmarks Introduction
  • CNN/DailyMail(Nallapati et al., 2016): A summarization dataset of news articles, with 500 randomly sampled items. (18–165 words)
  • HANNA(Chhun et al., 2022): A long-form story generation dataset with 200 selected items. (139–995 words)
  • TruthfulQA(Lin et al., 2021): A benchmark for factual accuracy in open-domain QA. (101–294 words)
  • HelloBench(Que et al., 2024): A long-text generation benchmark. We selected subsets from heuristic text generation (e.g., argumentative and roleplaying writing, covering five types) and open-ended QA (spanning ten domains). (489–1450 words)
  • GAOKAO-Bench(Zhang et al., 2023b): A benchmark collected from the Chinese college entrance examination (GAOKAO). We selected the 2010-2022 History Open-ended Questions subset. (71–901 words)
@@ Appendix E Detailed Experimental Results
Precise Length Constraint Benchmark
@@
Table 7: Combined Benchmark Evaluation Table
@@ E.2 Length Bias Correction in LLMs-as-a-Judge
E.3 Residual Length Error Analysis in MarkerGen
@@
E.4 Cross-layer Attention Analysis from the MarkerGen Perspective
@@
Combining the analyses from Figures 6 and 10, we infer that in the shallow layers, attention is primarily focused on the length information represented by the length markers. This suggests that the model’s early stages prioritize processing and understanding the input length. The higher entropy in these layers indicates that the model needs to integrate various details and information to effectively comprehend the input. As the model progresses to deeper layers, attention shifts from the length information to the adjacent semantic content. The lower entropy in these layers indicates that the model refines its focus, extracting key features and generating more relevant output.

### v3

@@ 1 Introduction
However, the ongoing enhancements in LLM capabilities have yet to deliver the expected performance in LCTG while ensuring semantic integrity Foster et al. (2024); Wang et al. (2024); Song et al. (2024). Yuan et al. (2024) reports that even advanced LLMs (e.g., GPT-4 Turbo (OpenAI, 2023)) violate the given length constraints almost 50% of the time. To address this, training-based methods (Park et al., 2024; Yuan et al., 2024; Jie et al., 2023; Li et al., 2024b) have been studied to reinforce LLMs’ adherence to length constraints, yet they face two key challenges: (1) Limited generalization: Since text types are diverse and length constraints vary widely (e.g., ranging from an exact 500 words to coarse intervals like 500-600 words or below 500 words), training-based methods often fail to generalize effectively across different settings, as demonstrated in Appendix E.1. (2) Inferior controllability: These methods strengthen LCTG by enforcing implicit length modeling during generation in a top-down manner via training, lacking the decomposition and targeted enhancement of LCTG sub-capabilities, thereby limiting their progress (Retkowski and Waibel, 2024).
@@ 2.5 LCTG Error
LCTG error is primarily attributed to fundamental deficiencies in length modeling, following the order of Counting Error $>$ Identifying Error $>$ Aligning Error $\gg$ Planning Error. Thus, as counting interval increases, the accumulation of counting errors leads to a corresponding rise in LCTG error.
@@ 3.2 Three-Stage Decoupled Generation
See Appendix  F for prompts of each stage.
@@ 4.3 Generalizability
Table 3 also shows MarkerGen’s strong performance across benchmarks with varying length scale (18-1450). To further investigate, we analyze progressively increasing the target length from 100 to 400. The results in Table 4 show a declining trend in MarkerGen’s error rate, which can be attributed to the auxiliary marker insertion decoding mechanism that prevents error accumulation from implicit modeling.
@@
[…]
@@ 4.4 Ablation Studies
The introduction of explicit length markers in the two-stage scheme leads to a substantial reduction in LCTG error relative to the implicit baseline $(8.7\rightarrow 2.66)$. However, this scheme places greater emphasis on length modeling, which consequently diminishes text quality $(4.45\rightarrow 4.28)$. In comparison, the three-stage scheme achieves a better balance by decoupling semantic and length modeling, thereby improving both length control and text quality.
@@ Conclusions
Acknowledgments
This work is supported by Beijing Natural Science Foundation (No.4222037, L181010).
@@ Appendix D Detailed Benchmarks Introduction
  • CNN/DailyMail(Nallapati et al., 2016): A summarization dataset of news articles, with 500 randomly sampled items. (18-165 words)
  • HANNA(Chhun et al., 2022): A long-form story generation dataset with 200 selected items. (139-995 words)
  • TruthfulQA(Lin et al., 2021): A benchmark for factual accuracy in open-domain QA. (101-294 words)
  • HelloBench(Que et al., 2024): A long-text generation benchmark. We selected subsets from heuristic text generation (e.g., argumentative and roleplaying writing, covering five types) and open-ended QA (spanning ten domains). (489-1450 words)
  • GAOKAO-Bench(Zhang et al., 2023b): A benchmark collected from the Chinese college entrance examination (GAOKAO). We selected the 2010-2022 History Open-ended Questions subset. (71-901 words)
@@ Appendix E Detailed Experimental Results
Precise Length Constraint Benchmarks
@@
Table 7: Evaluation of the training-based method Ruler on its own test set and four generalization benchmarks.
@@ E.1 Performance and Generalization Study of Training-based Methods
The test set of Ruler adopts two custom evaluation metrics: Precise Match (PM) and Flexible Match (FM). PM requires the output length to fall within a narrow target interval, while FM allows a broader tolerance range. The metrics are defined as:
$PM=\frac{1}{N}\sum_{i=1}^{N}1\left(\text{lb}^{P}_{\text{TL}_{i}}<L(c_{i})\leq\text{ub}^{P}_{\text{TL}_{i}}\right)$
$FM=\frac{1}{N}\sum_{i=1}^{N}1\left(\text{lb}^{F}_{\text{TL}_{i}}<L(c_{i})\leq\text{ub}^{F}_{\text{TL}_{i}}\right)$
where $L(c_{i})$ is the length of the $i$-th generated output, and $\text{TL}_{i}$ denotes the corresponding target length.
@@ E.2 Length Bias Correction in LLMs-as-a-Judge
E.3 Win-rate and LLM-as-Judge Based Evaluation
To provide a more robust and trustworthy evaluation of generation quality, we complement our main evaluation protocol with a win-rate-based assessment and multi-LLM judgment analysis. Specifically, we adopt the pairwise evaluation framework from AlpacaEval 2.0 Dubois et al. (2024), using GPT-4o-mini as the automatic judge. The generated outputs from the Implicit baseline and our proposed MarkerGen method are compared against reference model Gemini-2.0 FlashGoogle (2024) across four benchmarks. As shown in Table 9, MarkerGen consistently achieves higher win rates across most configurations, verifying the effectiveness of the Three-Stage Decoupled Generation in improving response quality.
To further confirm the reliability of using LLMs as judges, we conducted a human agreement study. We randomly sampled 200 pairwise comparisons and observed a strong Spearman’s correlation coefficient of 0.8794 between GPT-4o-mini judgments and human annotations, indicating high consistency.
In addition, we report point-wise quality scores evaluated by three different LLMs: GPT-4o, Gemini-2.0-Flash (thinking-exp), and GPT-4o-mini. Tables 101112, and 13 summarize these results across the CNN/DailyMail, HANNA, TruthfulQA, and Heuristic Generation benchmarks, respectively. Each table reports the average score ($S_{\text{avg}}$), variance ($\sigma^{2}$), and model sizes. MarkerGen consistently improves average scores across all model sizes and datasets, while maintaining comparable or lower variance, further reinforcing its robustness and generalizability.
Model Series Size Method CNN/DailyMail HANNA TruthfulQA Heuristic Gen.
Qwen2.5 32B Implicit 15.80% 6.50% 28.25% 1.63%
MarkerGen 16.43% 23.00% 36.91% 5.69%
14B Implicit 9.40% 7.04% 28.87% 1.63%
MarkerGen 11.60% 11.00% 24.12% 3.25%
7B Implicit 4.20% 1.51% 23.24% 0.00%
MarkerGen 14.80% 16.67% 25.62% 4.07%
Llama3.1 70B Implicit 7.40% 2.50% 33.81% 0.00%
MarkerGen 9.02% 6.03% 30.31% 1.63%
8B Implicit 12.27% 3.02% 19.90% 0.00%
MarkerGen 10.69% 4.15% 20.52% 0.00%
Table 9: Win rates (%) of Implicit baseline and MarkerGen compared against Gemini-2.0 Flash using GPT-4o-mini as judge across four benchmarks. Higher win rates indicate better generation quality.
Model Series Size Method S_4o S_gemini S_4omini S_avg $\sigma^{2}$
Qwen2.5 32B Implicit 3.02 3.64 3.21 3.29 0.067
MarkerGen 3.05 3.67 3.25 3.32 0.066
14B Implicit 2.92 3.55 3.15 3.21 0.099
MarkerGen 2.93 3.58 3.16 3.22 0.092
7B Implicit 2.68 3.39 3.04 3.04 0.125
MarkerGen 2.83 3.45 3.07 3.12 0.092
Llama3.1 70B Implicit 2.84 3.55 3.09 3.16 0.109
MarkerGen 3.08 3.69 3.36 3.38 0.068
8B Implicit 2.81 3.37 3.04 3.07 0.067
MarkerGen 2.78 3.38 3.18 3.11 0.074
Table 10: Evaluation scores of MarkerGen and Implicit baseline on CNN/DailyMail benchmark using three judges (GPT-4o, Gemini-2.0 Flash, GPT-4o-mini).
Model Series Size Method S_4o S_gemini S_4omini S_avg $\sigma^{2}$
Qwen2.5 32B Implicit 3.20 3.73 3.67 3.53 0.056
MarkerGen 3.29 3.75 3.72 3.59 0.044
14B Implicit 3.15 3.61 3.55 3.44 0.041
MarkerGen 3.16 3.61 3.55 3.44 0.039
7B Implicit 3.09 3.31 3.47 3.29 0.024
MarkerGen 3.12 3.36 3.50 3.33 0.024
Llama3.1 70B Implicit 3.18 3.66 3.61 3.48 0.047
MarkerGen 3.25 3.74 3.63 3.54 0.044
8B Implicit 3.13 3.45 3.54 3.37 0.030
MarkerGen 3.19 3.50 3.60 3.43 0.031
Table 11: Evaluation scores of MarkerGen and Implicit baseline on HANNA benchmark using three judges (GPT-4o, Gemini-2.0 Flash, GPT-4o-mini).
Model Series Size Method S_4o S_gemini S_4omini S_avg $\sigma^{2}$
Qwen2.5 32B Implicit 4.56 3.94 4.45 4.32 0.073
MarkerGen 4.63 4.09 4.54 4.42 0.055
14B Implicit 4.47 3.95 4.44 4.29 0.056
MarkerGen 4.44 3.93 4.43 4.27 0.058
7B Implicit 4.24 3.72 4.29 4.09 0.066
MarkerGen 4.34 3.83 4.33 4.17 0.056
Llama3.1 70B Implicit 4.53 4.09 4.46 4.36 0.038
MarkerGen 4.59 4.14 4.48 4.40 0.038
8B Implicit 4.20 3.62 4.22 4.01 0.076
MarkerGen 4.19 3.62 4.25 4.02 0.080
Table 12: Evaluation scores of MarkerGen and Implicit baseline on TruthfulQA benchmark using three judges (GPT-4o, Gemini-2.0 Flash, GPT-4o-mini).
Model Series Size Method S_4o S_gemini S_4omini S_avg $\sigma^{2}$
Qwen2.5 32B Implicit 4.53 4.14 3.80 4.16 0.089
MarkerGen 4.64 4.22 4.14 4.33 0.049
14B Implicit 4.42 3.98 3.80 4.07 0.068
MarkerGen 4.46 4.07 4.00 4.18 0.040
7B Implicit 3.97 3.70 3.42 3.69 0.050
MarkerGen 4.52 4.15 4.13 4.26 0.031
Llama3.1 70B Implicit 4.47 3.98 3.74 4.07 0.093
MarkerGen 4.54 4.08 3.98 4.20 0.058
8B Implicit 4.21 3.89 3.72 3.94 0.041
MarkerGen 4.47 4.04 4.03 4.18 0.042
Table 13: Evaluation scores of MarkerGen and Implicit baseline on Heuristic Generation benchmark using three judges (GPT-4o, Gemini-2.0 Flash, GPT-4o-mini).
E.4 Enhanced Ablation Study
To provide a more comprehensive validation of our method, we extend the original component ablation study (Table 5) by evaluating additional combinations of stages. This enhanced analysis is motivated by the need to isolate and quantify the individual and joint effects of the key stages, as suggested in the review. Table 14 presents results in terms of error score ($E\downarrow$), LLM-evaluated quality score ($S\uparrow$), and cost (relative to baseline, $\delta$).
Variant Components Interval $E$ ($\downarrow$) $S$ ($\uparrow$) Cost
Baseline (0) Direct Generation 64 30.27 4.51 $\delta$
w/o Tool (1) Plan + Generation 64 27.58 4.60 $1.05\delta$
Decaying 20.44 4.58 $1.56\delta$
w Tool (2) Insertion Strategy 64 5.96 4.27 $0.79\delta$
Decaying 3.58 4.27 $0.69\delta$
w/o Tool (w decouple) (3) Decoupled Generation 64 26.11 4.53 $1.20\delta$
Decaying 22.87 4.52 $1.11\delta$
Two Stage (1+2) Plan + Insertion Strategy 64 3.93 4.32 $0.84\delta$
Decaying 2.66 4.28 $0.79\delta$
w/o Tool (w plan & decouple) (1+3) Plan + Decoupled Generation 64 25.89 4.63 $1.56\delta$
Decaying 23.19 4.60 $1.51\delta$
w Tool (w/o plan & w decouple) (2+3) Insertion + Decoupled Generation Decaying 5.11 4.50 $1.10\delta$
Three Stage (1+2+3) Plan + Insertion + Decoupled Gen. Decaying 4.48 4.54 $1.46\delta$
Table 14: Extended ablation study analyzing the impact of different components combinations. $\delta$: cost relative to baseline.
As shown in Table 14, the removal of any individual stage leads to a significant increase in error or a drop in quality. Notably, the two-stage variant (1+2) achieves the lowest error (2.66) with a relatively low cost ($0.79\delta$), confirming its effectiveness and efficiency. Meanwhile, the full three-stage setup offers a balanced trade-off, delivering strong performance with moderate cost. These results reinforce the necessity of each stage and validate our full method design.
E.5 Residual Length Error Analysis in MarkerGen
@@
E.6 Cross-layer Attention Analysis from the MarkerGen Perspective
@@
Combining the analyses from Figures 6, 10, 11, 12, and 13, we infer that in the shallow layers, attention is primarily focused on the length information represented by the length markers. This suggests that the model’s early stages prioritize processing and understanding the input length. The higher entropy in these layers indicates that the model needs to integrate various details and information to effectively comprehend the input. As the model progresses to deeper layers, attention shifts from the length information to the adjacent semantic content. The lower entropy in these layers indicates that the model refines its focus, extracting key features and generating more relevant output.
@@
Appendix F Prompt for Three-Stage Decoupled Generation
The following three prompts correspond to the three stages of our proposed method. We present here the actual prompt templates used in the TruthfulQA task as representative examples: Stage One for planning (Figure 14), Stage Two for semantic integrity (Figure 15), and Stage Three for length-constrained rewriting (Figure 16).
Stage One Prompt: Planning Length Definition. A word is defined as any standalone word, number, or symbol, including punctuation and special symbols. You are tasked with generating a high-quality and truthful answer to the following question, aiming for approximately {target_length} words. Your answer must be factually accurate, free from any false information or hallucinations. Ensure that all statements can be supported by reliable sources. Planning Task: Understand the question: Comprehend what is being asked. Research and gather facts: Collect accurate and relevant information. Organize the response: Structure the answer logically. Ensure completeness: Address all aspects of the question. Maintain accuracy and relevance: Avoid unnecessary digressions. After planning, generate the answer based on this structure, maintaining logical consistency and factual accuracy.
Figure 14: Stage One Prompt: Planning
Stage Two Prompt: Semantic Integrity Answer generation task: Generate a comprehensive and precise answer to the question based on the plan, ensuring clarity, coherence, and factual accuracy. The answer should be approximately {target_length} words in length. A word is defined as any standalone word, number, or symbol, including punctuation and special symbols. Only the answer text should be output; do not add any extra comments, notes, or explanations. The answer should start with “Answer generation task:” and follow the format specified below. Place “###end” at the absolute end of the answer to mark its completion. Question: ‘‘{prompt}’’
Figure 15: Stage Two Prompt: Ensuring Semantic Integrity
Stage Three Prompt: Length-Constrained Rewriting Task Description: 1. In the first stage, we generated a high-quality answer without strict length control. 2. In this second stage, your task is to rewrite the high-quality answer to meet the specified length constraints. Rewriting Requirements: Preserve core meaning, accuracy, and factual correctness. Match the target length of {target_length} words as closely as possible. Shorten or expand while maintaining clarity and integrity. Insert or remove detail as needed without altering facts. Output only the answer; no commentary or explanation. Length Definition: A word is defined as any standalone word, number, or symbol, including punctuation and special symbols. Length Feedback: The high-quality answer contains {actual_length} words. It exceeds or falls short of the target length by {abs(length_difference)} words. Answer generation task: After planning the adjustments, rewrite the answer in one go, adhering to the planned structure and word count. Do not truncate unfinished sentences just to match the target. Insert length markers during generation: Start with larger intervals, then reduce spacing for detailed content. Markers should be numbered and evenly placed. Example 1 (Target length: 138 words): Answer generation task: The rain began to fall softly as the train sped through the countryside, its windows fogging up as the cool air met the warmth inside. The landscape outside blurred into a wash of green and gray as the train left the city behind, heading toward the mountains. Inside, the passengers sat quietly, some lost in their books, [64 words] others staring out at the passing scenery, their faces illuminated by the soft glow of the overhead lights. The rhythmic clattering of the train wheels on the tracks created a [96 words] calming rhythm, almost like a lullaby. As the train continued its journey, the fields [112 words] turned to forests, and the rivers widened [120 words]. A sense of peace settled over [128 words] the passengers as they [132 words] journeyed farther [134 words] into the [136 words] unknown [137 words]. [138 words] ###end Example 2 (Target length: 70 words): Answer generation task: The golden light of the setting sun bathed the city streets in a warm glow, [16 words] casting long shadows as people rushed home after a busy day. The streets buzzed with [32 words] activity, cars honking, and pedestrians chatting. Amid the hustle, a young couple [48 words] walked hand in hand, lost in conversation [56 words]. The sound of [60 words] their laughter mingled with [64 words] the noise [66 words] of the [68 words] city [69 words]. [70 words] ###end Final Prompt Input: {task_description}
{rewrite_requirements}
{length_definition}
The Question is {prompt}
First stage High-quality answer: {generated_answer}
{length_feedback}
{answer generation task}
Figure 16: Stage Three Prompt: Length-Constrained Rewriting